# Exceptional Tannaka groups only arise from cubic threefolds

### v2

@@ 1. Introduction
Any subvariety of an abelian variety gives rise to a reductive group via the convolution of perverse sheaves. For smooth subvarieties these Tannaka groups have recently been used to obtain arithmetic finiteness results for varieties over number fields [LS20, KM23] and the big monodromy criterion in [JKLM23]. For both applications one has to exclude exceptional Tannaka groups. Such exceptional groups are known to occur for Fano surfaces of smooth cubic threefolds, which have Tannaka group $E_{6}$ by [Krä16]. We show that in the range relevant for the above applications, these Fano surfaces are the only examples with exceptional Tannaka groups; this leads to a considerable strengthening of [KM23, JKLM23]. The key idea is of a general nature and goes far beyond the application in this paper: We upgrade the Tannaka formalism from perverse sheaves to complex Hodge modules in the sense of Sabbah-Schnell [SS18]. In particular, we prove a comparison theorem between the Tannaka groups for perverse sheaves and for Hodge modules which implies that the Hodge decomposition is defined by a cocharacter of the perverse Tannaka group. This puts a strong restriction on the arising Tannaka groups that will be useful also for other applications.
@@ 1.1. Tannaka groups of subvarieties
as the pushforward of the constant sheaf under the closed immersion $i\colon X\hookrightarrow A$, shifted in cohomological degree $-d$ so that it becomes an object of the abelian category $\textup{Perv}(A,\mathbb{F})$ of perverse sheaves on $A$ as in [BBDG18]. The group law on the abelian variety induces a convolution product on perverse sheaves so that the convolution powers of $\delta_{X}$ generate a neutral Tannaka category $\langle\delta_{X}\rangle$, see for instance [JKLM23, sect. 3]. In what follows, we fix a fiber functor
@@
which we call the Tannaka group of the subvariety $X\subset A$. If $X$ is invariant under translation by a nonzero point of $A$, then it descends to the quotient of $A$ modulo the subgroup generated by that point. Hence, in what follows we assume that $X$ is nondivisible in the sense that it is not stable under the translation by any nonzero point of $A$. In this case, the connected component of the identity $G_{X,\omega}^{\circ}\subset G_{X,\omega}$ acts irreducibly on the representation $\omega(\delta_{X})$ by [JKLM23, prop. 3.3]. We are interested in its derived group
@@
which is a connected semisimple algebraic group. The smoothness of $X$ implies that $\omega(\delta_{X})$ is a minuscule representation [Krä22, cor. 1.10] [JKLM23, cor 5.15] of dimension equal to the absolute value of the topological Euler caracteristic of $X$:
@@
This leaves only very few possibilities because the Tannaka group is known to be simple in most cases: Suppose that $g\geqslant 3$ and that the normal bundle of $X\subset A$ is ample. Then by [JKLM23, th. A], the group $G_{X,\omega}^{\ast}$ fails to be simple if and only if there exist smooth positive-dimensional subvarieties $X_{1},X_{2}\subset A$ such that the sum morphism induces an isomorphism
@@
Beyond the standard representations of classical groups, there are only few other minuscule representations of simple algebraic groups: Wedge powers of the standard representation of $\operatorname{SL}_{n}$, (half)spin representations of spin groups, and the smallest irreducible representations of the exceptional groups $E_{6}$ and $E_{7}$. In [JKLM23] we completely characterized wedge powers and ruled out spin representations; the main goal of this paper is to understand when exceptional groups occur.
@@ 1.2. Exceptional Tannaka groups
where $V$ is one of the two irreducible representations of $E_{6}$ with $\dim V=27$. It is clear from the description of $X$ as a Fano variety of lines that the classical Gauss map
@@
is an embedding, hence $X\subset A$ is nondivisible. Recall that $A$ is principally polarized and its theta divisor is the image of the difference morphism
@@
by [CG72, th. 13.4] [Bea82], which permits to recover the cubic threefold: The theta divisor has an isolated singularity at the origin and its tangent cone $C$ there can be identified with the cubic threefold $Y$ in suitable linear coordinates on $\operatorname{Lie}A$. In the same coordinates $Y$ is also identified with the image of the projection
@@
The goal of this paper is to show that under mild assumptions, the above is the only example of exceptional Tannaka groups. In particular, in the dimension range relevant for [KM23, JKLM23] we obtain (see corollary 7.2):
@@
Let $X\subset A$ be a smooth irreducible subvariety with ample normal bundle and dimension $<g/2$. Then the following are equivalent:
@@
The Fano surface $X$ of lines on a smooth cubic threefold $Y\subset\mathbb{P}^{4}$ does not always have ample normal bundle in the Albanese variety: By proposition 6.14 the normal bundle of an Albanese embedding $X\hookrightarrow\operatorname{Alb}(X)$ fails to be ample if and only if there is a hyperplane $H\subset\mathbb{P}^{4}$ such that the cubic surface $H\cap Y\subset H$ is a cone over an elliptic curve; this happens for instance for the cubic threefold with affine equation
[…]
where $f(x)$ is a degree $3$ polynomial without multiple roots. But $X$ still has a weaker positivity property: As in [JKLM23], we say that an integral subvariety $Z\subset A$ of an abelian variety $A$ is nondegenerate if for every surjective morphism of abelian varieties $\varphi\colon A\to B$ we have
@@
The Fano surface of lines on any smooth cubic threefold is nondegenerate inside its Albanese variety by [Sch17, th. 1], since it is a summand of a theta divisor; in fact loc. cit. refers to a stronger notion of nondegeneracy, which implies the above one by [Ran81, lemma II.12]. Our proof of A proceeds by reduction to the case of surfaces. In this case we only need to assume nondegeneracy, hence we obtain the following result which now includes the Fano surfaces of all smooth cubic threefolds (see theorem 7.1):
@@
    and the difference morphism $X\times X\to X-X$ has generic degree $6$.
@@
A different characterization has been obtained by Casalaina-Martin, Popa and Schreieder. In [CMPS18, th. 6.1] they show that if $X$ is a subvariety of minimal cohomology class of a principally polarized abelian fivefold $A$ and $X-X$ is a theta divisor, then $X$ is the Fano surface of lines on a smooth cubic threefold. Note that in both theorems A and B the canonical morphism $\operatorname{Alb}(X)\to X$ need not be an isomorphism:
@@
maps $X\subset\operatorname{Alb}(X)$ isomorphically onto its image $\bar{X}:=p(X)\subset A$. So this image is still an integral smooth nondivisible subvariety. The pushforward of perverse sheaves under an isogeny does not change the derived connected component of their Tannaka groups [JKLM23, cor. 3.5], so we still have
@@ 1.3. The Hodge filtration
A key ingredient in our proof of theorems A and B is the theory of complex Hodge modules by Sabbah-Schnell [SS18]. To explain the idea, let $A$ be a complex abelian variety and $\operatorname{HM}(A)$ the category of direct sums of complex Hodge modules of arbitrary weight on it, see section 2 for details. The category $\operatorname{HM}(A)$ is semisimple $\mathbb{C}$-linear abelian and comes equipped with a faithful exact $\mathbb{C}$-linear functor
@@
To any $M\in\operatorname{HM}(A)$ one may attach a Tannaka group $G_{\omega}(M)$ in the same way as for perverse sheaves; in corollary 2.5 we show a comparison theorem between this group and the Tannaka group $G_{\omega}(P)$ of the perverse sheaf $P=\operatorname{DR}(M)$. In particular, the derived groups of their connected components of the identity are the same:
@@
This allows us to enrich with Hodge theoretic data all the previous Tannakian constructions for perverse sheaves, so for the rest of this section we will work only with Hodge modules. It is now time to be more specific about our choice of fiber functors: Fix a Hodge module $M\in\operatorname{HM}(A)$, and let $L$ be a unitary local system of rank one on $A$ with the property that all perverse subquotients of $\operatorname{DR}(M)\otimes L$ have their cohomology concentrated in degree zero; there are plenty of such local systems by generic vanishing for perverse sheaves on abelian varieties [KW15, Sch15]. Then the functor
@@
We show that this Hodge decomposition is compatible with the convolution product, which allows to enrich $\omega$ to a tensor functor $\omega^{H}\colon\langle M\rangle\rightarrow\operatorname{Vect}_{\mathbb{Z}\times\mathbb{Z}}(\mathbb{C})$ with values in the category of finite-dimensional bigraded complex vector spaces. This leads to the following result (see corollary 3.2):
@@
see section 3.2. Comparing this with an improvement to the Hodge estimates by Lazarsfeld-Popa [LP10] and Lombardi [Lom13] explained in appendix A, we obtain the following result (see proposition 4.6):
@@
    $G_{X,\omega}^{\ast}\simeq E_{6}$$|\chi_{\textup{top}}(X)|=27$ and $d\in\{2,4,6\}$, or
@@
    $G_{X,\omega}^{\ast}\simeq E_{7}$$|\chi_{\textup{top}}(X)|=56$ and $d\in\{3,\dots,15\}$ is odd.
@@
In each case we also know the Hodge numbers $h^{p}(X):=(-1)^{d-p}\chi(X,\Omega^{p}_{X})$, as they must be among those listed in propositions 4.3, LABEL: and 4.5. This essentially reduces the proof of theorem A to the case of surfaces treated in theorem B. For the proof of the latter we apply the representation theory of $E_{6}$ to show that the difference morphism
@@
is generically finite of degree $\geqslant 6$ over its image (see section 4.6), and then use this numerical information to show by direct geometric arguments that the projective tangent cone to $D\subset A$ at the origin is a smooth cubic threefold whose associated Fano surface of lines is isomorphic to $X$ (see section 7).
@@ 1.4. No country for $E_{7}$
Although D imposes strong restrictions, it is not enough to rule out completely the existence of subvarieties $X$ with $G_{X,\omega}^{\ast}\simeq E_{7}$ in the dimension range $d<g/2$. With the $E_{6}$ case in mind, one might think that a natural candidate for such an $X$ would be the Fano variety of lines on a quartic double solid, that is, a double cover of $\mathbb{P}^{3}$ branched along a smooth quartic. Indeed, the $(-1)$-curves on a del Pezzo surface of degree $3$ resp. $2$ form a root system of type $E_{6}$ resp. $E_{7}$. A cubic threefold can be thought of as a one parameter family of del Pezzo surfaces of degree $3$ since these are cubic surfaces. Similarly, a quartic double solid is a one parameter family of del Pezzo surfaces of degree $2$ since the anticanonical divisor realizes them as double covers of $\mathbb{P}^{2}$ branched along smooth quartic curves.
However, the pieces of the jigsaw puzzle do not quite match together in the $E_{7}$ case: For instance, lines on a quartic double solid are parametrized by a surface, while D states that the sought-for example must be a threefold. This parity issue is hardwired in the representation theory of $E_{7}$ and lies at the heart of the following negative result (see corollary 8.2):
@@
For any smooth irreducible subvariety $X\subset A$ with ample normal bundle and dimension $<g/2$ we have
@@
The proof of this will be given in section 8. It does not require Hodge modules and is largely independent from the rest of this paper: The idea is to control the decomposition of the tensor square $\delta_{X}*\delta_{X}$ by looking at the fibers of the sum morphism $X\times X\to A$ and using Kashiwara’s estimate for characteristic varieties of direct images. Once we know how the tensor square decomposes, we can then conclude by a version of Larsen’s alternative.
@@ 1.5. Big Monodromy
As an application of the above results we can significantly strengthen the big monodromy criterion in [JKLM23]. More precisely, let $S$ be a smooth complex variety. Let $A$ be a complex abelian variety of dimension $g$ and $\mathcal{X}\subset A_{S}:=A\times S$ a subvariety such that the projection $f\colon\mathcal{X}\to S$ is a smooth morphism with connected fibers of dimension $d$. The situation is resumed in the following diagram:
@@
The fiber at $s$ of this local system comes with a linear action of the group $\pi_{1}(S,s)$ via the monodromy representation. With terminology as in loc. cit. we obtain the following strengthening of the big monodromy result in [JKLM23, sect. 1.1]:
@@
Suppose $X:=\mathcal{X}_{\bar{\eta}}\subset A_{S,\bar{\eta}}$ has ample normal bundle and dimension $d<(g-1)/2$. If $\mathcal{X}$ is symmetric up to translation, assume that (1.1) below holds. Then the following are equivalent:
@@
(1.1) $|e|\;\neq\;2^{2m-1}\quad\text{if\leavevmode\nobreak\ $d\geqslant(g-1)/4$ and $m\in\{3,\dots,d\},m\equiv d$ modulo $2$.}$
This assumption is empty for $d<(g-1)/4$. Apart from curves, in [JKLM23] we had no control on the dimension and irregularity of varieties with exceptional Tannaka group, so we assumed $|e|\neq 27,56$ for all $(d,g)$ in the given range. Theorems A and E now permit to remove these assumptions completely. For applications outside this range, we only need to exclude the finite list of $(d,g)$ in D.

### v3

@@ Abstract.
Key words and phrases:
Subvarieties of abelian varieties, perverse sheaves, Hodge modules, Tannaka categories, exceptional groups, Fano surfaces, cubic threefolds.
2020 Mathematics Subject Classification:
14K12 (primary), 14D05, 14J70, 18M25, 20G05, 32S60 (secondary).
@@ 1. Introduction
Any subvariety of an abelian variety gives rise to a reductive group via the convolution of perverse sheaves. For smooth subvarieties these Tannaka groups have recently been used to obtain arithmetic finiteness results for varieties over number fields [LS20, KM23] and the big monodromy criterion in [JKLM25]. For both applications one has to exclude exceptional Tannaka groups. Such exceptional groups are known to occur for Fano surfaces of smooth cubic threefolds, which have Tannaka group $E_{6}$ by [Krä16]. We show that in the range relevant for the above applications, these Fano surfaces are the only examples with exceptional Tannaka groups; this leads to a considerable strengthening of [KM23, JKLM25]. The key idea is of a general nature and goes far beyond the application in this paper: We upgrade the Tannaka formalism from perverse sheaves to complex Hodge modules in the sense of Sabbah-Schnell [SS18]. In particular, we prove a comparison theorem between the Tannaka groups for perverse sheaves and for Hodge modules which implies that the Hodge decomposition is defined by a cocharacter of the perverse Tannaka group. This puts a strong restriction on the arising Tannaka groups that will be useful also for other applications.
@@ 1.1. Tannaka groups of subvarieties
as the pushforward of the constant sheaf under the closed immersion $i\colon X\hookrightarrow A$, shifted in cohomological degree d-d so that it becomes an object of the abelian category $\textup{Perv}(A,\mathbb{F})$ of perverse sheaves on $A$ as in [BBDG18]. The group law on the abelian variety induces a convolution product on perverse sheaves so that the convolution powers of $\delta_{X}$ generate a neutral Tannaka category $\langle\delta_{X}\rangle$, see for instance [JKLM25, sect. 3]. In what follows, we fix a fiber functor
@@
which we call the Tannaka group of the subvariety $X\subset A$. If $X$ is invariant under translation by a nonzero point of $A$, then it descends to the quotient of $A$ modulo the subgroup generated by that point. Hence, in what follows we assume that $X$ is nondivisible in the sense that it is not stable under the translation by any nonzero point of $A$. In this case, the connected component of the identity $G_{X,\omega}^{\circ}\subset G_{X,\omega}$ acts irreducibly on the representation $\omega(\delta_{X})$ by [JKLM25, prop. 3.3]. We are interested in its derived group
@@
which is a connected semisimple algebraic group. The smoothness of $X$ implies that $\omega(\delta_{X})$ is a minuscule representation [Krä22, cor. 1.10] [JKLM25, cor 5.15] of dimension equal to the absolute value of the topological Euler characteristic of $X$:
@@
This leaves only very few possibilities because the Tannaka group is known to be simple in most cases: Suppose that $g\geqslant 3$ and that the normal bundle of $X\subset A$ is ample. Then by [JKLM25, th. A], the group $G_{X,\omega}^{\ast}$ fails to be simple if and only if there exist smooth positive-dimensional subvarieties $X_{1},X_{2}\subset A$ such that the sum morphism induces an isomorphism
@@
Beyond the standard representations of classical groups, there are only few other minuscule representations of simple algebraic groups: Wedge powers of the standard representation of $\operatorname{SL}_{n}$, (half)spin representations of spin groups, and the smallest irreducible representations of the exceptional groups $E_{6}$ and $E_{7}$. In [JKLM25] we completely characterized wedge powers and ruled out spin representations; the main goal of this paper is to understand when exceptional groups occur.
@@ 1.2. Exceptional Tannaka groups
where $V$ is one of the two irreducible representations of $E_{6}$ with $\dim V=27$. Via the Tannaka formalism, the trilinear form defining $E_{6}$ corresponds to an irreducible fiber of the sum map $X\times X\times X\to A$ of dimension $3$ parametrizing triples of coplanar lines in $Y$. It is clear from the description of $X$ as a Fano variety of lines that the classical Gauss map
@@
is an embedding, hence $X\subset A$ is nondivisible. The difference morphism
@@
is generically finite of degree $6$ and its fibers are given by classical “double-six” configurations of lines on smooth cubic surfaces [CG72, sect. 13]. Its image is a theta divisor of the principally polarized abelian variety $A$ and permits to recover the cubic threefold [CG72, th. 13.4] [Bea82]: The theta divisor has an isolated singularity at the origin and its tangent cone $C$ there can be identified with the cubic threefold $Y$ in suitable linear coordinates on $\operatorname{Lie}A$. In the same coordinates $Y$ is also identified with the image of the projection
@@
The goal of this paper is to show that under mild assumptions, the above is the only example of exceptional Tannaka groups. In particular, in the dimension range relevant for [KM23, JKLM25] we obtain (see corollary˜7.2):
@@
Let $X\subset A$ be a smooth irreducible subvariety with ample normal bundle and dimension <g/2<g/2. Then the following are equivalent:
@@
The Fano surface $X$ of lines on a smooth cubic threefold $Y\subset\mathbb{P}^{4}$ does not always have ample normal bundle in its Albanese variety: This fails for example for the cubic threefold with affine equation
$y^{2}-f(x)+z^{2}t+t^{3}\;=\;0,$
where $f(x)$ is a degree $3$ polynomial without multiple roots. In fact, it is easy to see that the normal bundle of an Albanese embedding $X\hookrightarrow\operatorname{Alb}(X)$ fails to be ample if and only if there is a hyperplane $H\subset\mathbb{P}^{4}$ such that the cubic surface $H\cap Y\subset H$ is a cone over an elliptic curve. Such hyperplanes are in bijection with the vertices of the corresponding cones, which are the Eckardt points of $Y$. There are at most finitely many such Eckardt points and their absence is equivalent to ampleness of the cotangent bundle of $X$. See [Rou09, sect. 2] for details.
To include cubic threefolds with Eckardt points, we relax the ampleness of the normal bundle to a weaker positivity property: We say that an integral subvariety $Z\subset A$ of an abelian variety $A$ is nondegenerate if for every surjective morphism of abelian varieties $\varphi\colon A\to B$ we have
@@
The Fano surface of lines on any smooth cubic threefold is nondegenerate inside its Albanese variety by [Sch17, th. 1], since it is a summand of a theta divisor; in fact loc. cit. refers to a stronger notion of nondegeneracy, which implies the above one by [Ran81, lemma II.12]. Our proof of ˜A proceeds by reduction to the case of surfaces. In this case we only need to assume nondegeneracy, hence we obtain the following result which now includes the Fano surfaces of all smooth cubic threefolds (see theorem˜7.1):
@@
    the difference morphism $X\times X\to X-X$ has generic degree $\geqslant 6$, and the sum morphism $X\times X\times X\to A$ has an irreducible fiber of dimension $\geqslant 3$.
@@
A different characterization has been obtained by Casalaina-Martin, Popa and Schreieder in [CMPS18, th. 6.1]: They show that if $X$ is a subvariety of minimal cohomology class of a principally polarized abelian fivefold $A$ and XXX-X is a theta divisor, then $X$ is the Fano surface of lines on a smooth cubic threefold. Note that in both theorems A and B, the canonical morphism $\operatorname{Alb}(X)\to A$ need not be an isomorphism:
@@
maps $X\subset\operatorname{Alb}(X)$ isomorphically onto its image $\bar{X}:=p(X)\subset A$. So this image is still an integral smooth nondivisible subvariety. The pushforward of perverse sheaves under an isogeny does not change the derived connected component of their Tannaka groups [JKLM25, cor. 3.5], so we still have
@@ 1.3. The Hodge filtration
A key ingredient in our proof of theorems A and B is the theory of complex Hodge modules by Sabbah-Schnell [SS18]. To explain the idea, let $A$ be a complex abelian variety and $\operatorname{HM}(A)$ the category of direct sums of complex Hodge modules of arbitrary weight on it, see section˜2 for details. The category $\operatorname{HM}(A)$ is semisimple $\mathbb{C}$-linear abelian and comes equipped with a faithful exact $\mathbb{C}$-linear functor
@@
To any $M\in\operatorname{HM}(A)$ one may attach a Tannaka group $G_{\omega}(M)$ in the same way as for perverse sheaves; in corollary˜2.5 we show a comparison theorem between this group and the Tannaka group $G_{\omega}(P)$ of the perverse sheaf $P=\operatorname{DR}(M)$. In particular, the derived groups of their connected components of the identity are the same:
@@
This allows us to enrich with Hodge theoretic data all the previous Tannaka constructions for perverse sheaves, so for the rest of this section we will work only with Hodge modules. It is now time to be more specific about our choice of fiber functors: Fix a Hodge module $M\in\operatorname{HM}(A)$, and let $L$ be a unitary local system of rank one on $A$ with the property that all perverse subquotients of $\operatorname{DR}(M)\otimes L$ have their cohomology concentrated in degree zero; there are plenty of such local systems by generic vanishing for perverse sheaves on abelian varieties [KW15b, Sch15]. Then the functor
@@
We show that this Hodge decomposition is compatible with the convolution product, which allows to enrich $\omega$ to a tensor functor $\omega^{H}\colon\langle M\rangle\rightarrow\operatorname{Vect}_{\mathbb{Z}\times\mathbb{Z}}(\mathbb{C})$ with values in the category of finite-dimensional bigraded complex vector spaces. This leads to the following result (see corollary˜3.2):
@@
see section˜3.2. Comparing this with an improvement to the Hodge estimates by Lazarsfeld-Popa [LP10] and Lombardi [Lom13] explained in appendix˜A, we obtain the following result (see proposition˜4.6):
@@
    $G_{X,\omega}^{\ast}\simeq E_{6}$, $|\chi_{\textup{top}}(X)|=27$ and $d\in\{2,4,6\}$, or
@@
    $G_{X,\omega}^{\ast}\simeq E_{7}$, $|\chi_{\textup{top}}(X)|=56$ and $d\in\{3,\dots,15\}$ is odd.
@@
In each case we also know the Hodge numbers $h^{p}(X):=(-1)^{d-p}\chi(X,\Omega^{p}_{X})$, as they must be among those listed in propositions˜4.3, LABEL: and 4.5. This essentially reduces the proof of theorem A to the case of surfaces treated in theorem B. For the proof of the latter we apply the representation theory of $E_{6}$ to show that the difference morphism
@@
is generically finite of degree $\geqslant 6$ over its image (see section˜4.6), and then use this numerical information to show by direct geometric arguments that the projective tangent cone to $D\subset A$ at the origin is a smooth cubic threefold whose associated Fano surface of lines is isomorphic to $X$ (see section˜7).
@@ 1.4. No country for $E_{7}$
Although ˜D imposes strong restrictions, it is not enough to rule out completely the existence of subvarieties $X$ with $G_{X,\omega}^{\ast}\simeq E_{7}$ in the dimension range $d<g/2$. With the $E_{6}$ case in mind, one might think that a natural candidate for such an $X$ would be the Fano variety of lines on a quartic double solid, that is, a double cover of $\mathbb{P}^{3}$ branched along a smooth quartic. Indeed, the (1)(-1)-curves on a del Pezzo surface of degree $3$ resp. $2$ give rise to a root system of type $E_{6}$ resp. $E_{7}$. A cubic threefold can be thought of as a one parameter family of del Pezzo surfaces of degree $3$ since these are cubic surfaces. Similarly, a quartic double solid is a one parameter family of del Pezzo surfaces of degree $2$ since the anticanonical divisor realizes them as double covers of $\mathbb{P}^{2}$ branched along smooth quartic curves.
However, the pieces of the jigsaw puzzle do not quite match together in the $E_{7}$ case: For instance, lines on a quartic double solid are parametrized by a surface, while ˜D states that the sought-for example must be a threefold. This parity issue is hardwired in the representation theory of $E_{7}$ and lies at the heart of the following negative result (see corollary˜8.2):
@@
For any smooth irreducible subvariety $X\subset A$ with ample normal bundle and dimension <g/2<g/2 we have
@@
The proof of this will be given in section˜8. It does not require Hodge modules and is largely independent from the rest of this paper: The idea is to control the decomposition of the tensor square $\delta_{X}*\delta_{X}$ by looking at the fibers of the sum morphism $X\times X\to A$ and using Kashiwara’s estimate for characteristic varieties of direct images. Once we know how the tensor square decomposes, we can then conclude by a version of Larsen’s alternative.
@@ 1.5. Big Monodromy
As an application of the above results we can significantly strengthen the big monodromy criterion in [JKLM25]. More precisely, let $S$ be a smooth complex variety. Let $A$ be a complex abelian variety of dimension $g$ and $\mathcal{X}\subset A_{S}:=A\times S$ a subvariety such that the projection $f\colon\mathcal{X}\to S$ is a smooth morphism with connected fibers of dimension $d$. The situation is summarized in the following diagram:
@@
The fiber at $s$ of this local system comes with a linear action of the group $\pi_{1}(S,s)$ via the monodromy representation. With terminology as in loc. cit. we obtain the following strengthening of the big monodromy result in [JKLM25, sect. 1.1]:
@@
Suppose $X:=\mathcal{X}_{\bar{\eta}}\subset A_{S,\bar{\eta}}$ has ample normal bundle and dimension d<(g1)/2d<(g-1)/2. If $\mathcal{X}$ is symmetric up to translation, assume that (1.1) below holds. Then the following are equivalent:
@@
(1.1) $|e|\;\neq\;2^{2m-1}\quad\text{if\penalty 10000\ $d\geqslant(g-1)/4$ and $m\in\{3,\dots,d\},m\equiv d$ modulo $2$.}$
This assumption is empty for d<(g1)/4d<(g-1)/4. Apart from curves, in [JKLM25] we had no control on the dimension and irregularity of varieties with exceptional Tannaka group, so we assumed $|e|\neq 27,56$ for all $(d,g)$ in the given range. Theorems A and E now permit to remove these assumptions completely. For applications outside this range, we only need to exclude the finite list of $(d,g)$ in ˜D.